OpenStudy (dtan5457):

Can someone guide me through finding sin 105 degrees using the half angle formula?

OpenStudy (dtan5457):

This is what I have so far. 210 degrees \[\pm \sqrt{\frac{ 1-\frac{ \sqrt{3} }{ 2 } }{ 2 }}\]

OpenStudy (dtan5457):

so far so good?

OpenStudy (amorfide):

yes because you have subbed into a half angle formula for sinx/2

OpenStudy (dtan5457):

then next step, i should just to cancel some denominators by multipying each term by 2?

OpenStudy (amorfide):

I would get rid of the fraction in the top of the fraction first by multiplying top and bottom by 2 then trying to bring stuff out of the root

OpenStudy (superdavesuper):

hmmm try 210 = 180 + 30 and then apply sin(A+B) = sinAcosB + cosAsinB...

OpenStudy (amorfide):

that is addition formula not half angle formula is it not?

OpenStudy (amorfide):

anyway carry on, since you have done it correctly assuming that cos(105)=root3/2

OpenStudy (dtan5457):

so leave the denominator alone for now, and just multiply the top part by 2/2?

OpenStudy (amorfide):

\[\frac{ 1-\frac{ \sqrt{3} }{ 2} }{ 2 }= \frac{ 2(1)- \frac{ 2\sqrt{3} }{ 2 } }{2(2) }\]

OpenStudy (amorfide):

then obviously put your square root back

OpenStudy (amorfide):

after that step you are pretty much finished to be honest

OpenStudy (dtan5457):

im getting \[\frac{ \sqrt{2-\sqrt{3}} }{ 2 }\]

OpenStudy (dtan5457):

is my answer correct?

OpenStudy (amorfide):